A necklace is made of 24 beads that are either white or green. The ratio of white beads to green beads in this necklace is 5 to 1. A second necklace is made of 40 beads that are either white or green. The ratio of white beads to green beads in this second necklace is 5 to 3. How many more green beads are in the necklace with 40 beads than in the necklace with 24 beads?  

Answer by ikleyn(52852)

.
A necklace is made of 24 beads that are either white or green.
The ratio of white beads to green beads in this necklace is 5 to 1.
A second necklace is made of 40 beads that are either white or green.
The ratio of white beads to green beads in this second necklace is 5 to 3.
How many more green beads are in the necklace with 40 beads than in the necklace with 24 beads?
~~~~~~~~~~~~~~~~


                    I will solve it step by step.


(a)  24 beads necklace: 5x + x = 24;  6x = 24;  x = 24/6 = 4.

         There are x = 4 green beads.



(b)  40 beads necklace: 5y + 3y = 40;  8y = 40;  y = 40/8 = 5.

         There are 3y = 3*5 = 15 green beads.



(c)  Now, to answer the question, you need subtract 4 from 15.

Solved.